A bag contains 2 red marbles, 3 blue marbles, and 7b green marbles. One marble is randomly selected from the bag.Find the following probabilities: a) It was not a blue marble b)It was red or blue marble
Answer #1

a)

P(not blue) = 1 - P(Blue)

= 1 - n(blue) / n(S)

= 1 - [ 3 / (2 + 3 + 7)

= 0.75

b)

P(Red or blue) = P(Red) + P(Blue)

= n(red) / n(S) + n(Blue ) / n(S)

= 2 / 12 + 3 / 12

= 0.4167